OpenStudy (anonymous):

What is the angle of refraction when a ray of light passes from ethanol to air if the incident angle is 34°? The refractive index of ethanol is 1.36 and of air is 1.00.

OpenStudy (anonymous):

apply Snell's Law to solve this

OpenStudy (anonymous):

\[n_{1}\sin \theta_{1}=n_{2}\sin \theta_{2}\] \[1.36*\sin 34=1.00*\sin \theta_{2}\]

OpenStudy (anonymous):

Put value of sin34 in the LHS and then solve for getting value of angle on the right